# Right Place, Right Time! Dynamizing Topological Graphs for Embodied Navigation

### v3

@@ Abstract
Embodied Navigation tasks often involve constructing topological graphs of a scene during exploration to facilitate high-level planning and decision-making for execution in continuous environments. Prior literature makes the assumption of static graphs with stationary targets, which does not hold in many real-world environments with moving objects. To address this, we present a novel formulation generalizing navigation to dynamic environments by introducing structured object transitions to dynamize static topological graphs called Object Transition Graphs (OTGs). OTGs simulate portable targets following structured routes inspired by human habits. We apply this technique to Matterport3D (MP3D), a popular simulator for evaluating embodied tasks. On these dynamized OTGs, we establish a navigation benchmark by evaluating Oracle-based, Reinforcement Learning, and Large Language Model (LLM)-based approaches on a multi-object finding task. Further, we quantify agent adaptability, and make key inferences such as agents employing learned decision-making strategies generalize better than those relying on privileged oracle knowledge. To the best of our knowledge, ours is the first work to introduce structured temporal dynamism on topological graphs for studying generalist embodied navigation policies. The code and dataset for our OTGs will be made publicly available to foster research on embodied navigation in dynamic scenes.
Figure 1: Object Transition Graph (OTGs): We introduce dynamism to topological graphs via portable targets following structured object transition scenarios. In this representative figure, the watch (in red) moves from node N1 at T=10:00 AM to node N5 at T=7:45 PM, while the wallet (in orange) moves from node N5 at T=10:15 AM to node N1 at T=7:45 PM. We study the performance of various navigation agents in different object transition scenarios, and establish their generalizability.
@@ 1 Introduction
Embodied navigation tasks often rely on topological graphs (TGs) for decision-making. Visual-Language Navigation (VLN) tasks such as REVERIE [36] and R2R [2, 11] utilize TGs extracted from MP3D [8] to follow human guidance. Similarly, methods for ObjectNav, PointNav, and ImageNav often employ TGs  [20, 53, 14, 55, 10, 22, 38], where nodes represent locations or landmarks, and edges define traversable paths. These graphs provide a structured way to plan and execute navigation policies.
@@
A key limitation of these existing embodied navigation paradigms is their reliance on static graphs, where objects within nodes remain stationary. This static assumption is unrealistic, as real-world environments are inherently dynamic, where users frequently move small, portable objects from place to place, such as shifting their phone or wallet between rooms throughout the day. Studies on robots in collaborative environments  [34, 3, 16] suggest that human object placements follow structured patterns shaped by routines and habits, introducing a degree of entropy in object locations over time. Despite this, existing navigation benchmarks fail to model such temporal variability, limiting their applicability for real-world deployment.
@@
In the ObjectNav task [18, 4] for instance, common state-of-the-art navigation approaches involve Reinforcement Learning (RL) [54, 52, 4, 47] and LLM-based zero-shot methods [18, 45, 6] to find stationary objects. RL policies for multi-object settings [50, 9] have also been introduced, but these are also based on static environments. For real-world deployment, these approaches must generalize, or maintain performance in dynamic scenes with shifting targets. Generalizability in prior embodied work has been described in terms of measuring an agent’s capacity to adapt to novel settings, but this usually refers to static environments that were unseen or used synthesized guidance language [57, 26, 23, 17]. However, there has been little to no work that defines generalizability in the context of adapting to dynamic settings with moving objects. The evaluation of these prior approaches on dynamic settings with non-stationary objects remains to be seen.
To address this gap, in this work, we introduce Object-Transition Graphs (OTGs), a dynamic variant of topological graphs that incorporates structured object transitions over time. Figure 1 illustrates this concept. Unlike traditional TGs, where navigation is performed on a static graph, OTGs introduce a temporal component, where objects shift positions while the underlying graph structure remains constant. This forces an agent to synchronize its trajectory with temporal object movements; its adaptability to change allows us to measure its generalist navigation performance. In dynamizing static topological graphs, our work seeks to redefine how we evaluate decision-making on embodied agents, by emphasizing not just where to navigate but also when.
In this work, we define the generalizability of a navigation approach by comparing the performance between static TGs and their OTG counterparts. This allows us to quantify an agent’s robustness to dynamic changes, providing us with key insights on the effectiveness of graph-based navigation schemes in real-world settings.
Main Results. We generalize embodied navigation to dynamic environments by 1) introducing structured object transitions to make static topological graphs dynamic, and 2) presenting a benchmark for evaluating the adaptability of state-of-the-art navigation policies in such dynamic settings. Our contributions are summarized as follows:
@@
  • Dynamizing Topological Graphs: We propose a novel framework for introducing structured object transitions on static topological graphs, transforming them into Object Transition Graphs or OTGs. Inspired by human habits, our approach defines three distinct object transition scenarios to govern the movement of small, portable objects in a scene, providing varying levels of entropy. Our method applies to any static topological graph, including those obtained from Matterport3D (MP3D) [8] and HM3D [37]. Our dynamized MP3D environments and code will be released as an open-source benchmark to foster research in this area.
  • OTG Benchmark: We introduce a benchmark for evaluating navigation performance on OTGs, comparing heuristic, RL, and LLM-based agents across structured object transition scenarios on a multi-object finding task. Measuring performance of agents on OTGs provides insights into the adaptability of graph-based approaches in dynamic environments with non-stationary objects.
  • Generalization in Dynamic Environments: We make several key inferences on the generalizability of various navigation approaches by comparing their performance on OTGs against static TGs. In particular, we observe that strategies relying on privileged knowledge can lead to suboptimal decision-making in dynamic environments, while approaches that improve agents with experience and collected observations better generalize to OTGs.
@@ 2 Related Works
Figure 2: Object Transition: Portable objects move around the scene at various timesteps in accordance to their natural rooms (Table 1) and transit scenarios (Table 2). Here, a mug is placed on a kitchen node $N_{3}$ at timestep $T=53$, but moves to a bedroom node $N_{5}$ at timestep $T=55$ via node $N_{4}$. If an agent reaches the kitchen after $T=53$ (or the bedroom before $T=55$), it would fail to see the mug. Multiple objects could also be at the same node (hat and mug in $N4$).
Traditionally, navigation in dynamic environments has focused on low-level planning tasks, such as crowd avoidance [41, 31, 5] and socially aware navigation [7, 29], where agents continuously maneuver to avoid obstacles while minimizing trajectory length or time. In contrast, embodied navigation tasks employing high-level planning on graphs are predominantly studied in static environments, with limited research on handling dynamic targets  [11, 44]. With OTGs, we introduce a form of structured temporality in dynamic environments where agents must reason over spatio-temporal changes, extending high-level planning to dynamic scenes with shifting targets. While recent simulators like Habitat 3.0 [35] include object rearrangement tasks, these primarily focus on agents modifying the environment rather than adapting to dynamic object placements over time.
@@
Rudra et al. [40] define small portable objects around the house and propose a contextual bandit scheme that aims to learn the likelihood of finding an object at various waypoints. In their case, however, object locations are shuffled only after each episode, meaning it finally boils down to an object-finding navigation task in a static environment. In contrast, we tackle a truly dynamic case, where objects are moving even during the episode. This definition adds a layer of complexity as the embodied agent must now navigate towards a constantly shifting target object, for which it needs to identify specific routines and object movement patterns in the environment.
Kurenkov et al. [21] introduced dynamizing household environments and experimented with scene graph memory to predict object locations. They also performed target object-finding experiments, but the environment was static as the agent moved. In our work, the objects move as the agent moves, and we deal with topological graphs, not scene graphs. Furthermore, Wang et al. [49] used LLM-generated human activities to dynamize a topological graph of a household environment and performed ObjectNav experiments in these environments. In contrast, our work focuses on converting topological graphs into object transition graphs, and comments on the generalizability of embodied navigation approaches in dynamic environments.
3 Object Transition Graph (OTG):
We define Object Transition Graphs (OTGs) is an evolutionary undirected graph where the underlying graph structure remains constant, but the objects present in its nodes evolve over time. Formally, let $G=(V,E)$ be a topological graph structure with nodes $V$ representing spatial locations (rooms and objects), and edges $E$ representing traversable paths between the nodes. Let $\mathcal{O}$ represent the total set of available portable objects. Then, at each timestep $t$, $v\in V$ is associated with a set of portable target objects $O_{v}(t)\subseteq\mathcal{O}$. While the graph topology remains static, the object assignments $O_{v}(t)$ evolve via a time-dependent transitioning style $\Lambda$, to give us an object transition function $O$ as,

### v4

# TAS: A Transit-Aware Strategy for Embodied Navigation with Non-Stationary Targets

@@ Abstract
Embodied navigation methods commonly operate in static environments with stationary targets. In this work, we present a new algorithm for navigation in dynamic scenarios with non-stationary targets. Our novel Transit-Aware Strategy (TAS) enriches embodied navigation policies with object path information. TAS improves performance in non-stationary environments by rewarding agents for synchronizing their routes with target routes. To evaluate TAS, we further introduce Dynamic Object Maps (DOMs), a dynamic variant of node-attributed topological graphs with structured object transitions. DOMs are inspired by human habits to simulate realistic object routes on a graph. Our experiments show that on average, TAS improves agent Success Rate (SR) by 21.1 in non-stationary environments, while also generalizing better from static environments by 44.5% when measured by Relative Change in Success (RCS). We qualitatively investigate TAS-agent performance on DOMs and draw various inferences to help better model generalist navigation policies. To the best of our knowledge, ours is the first work that quantifies the adaptability of embodied navigation methods in non-stationary environments. Code and data for our benchmark will be made publicly available.
@@ 1 Introduction
Embodied navigation involves an agent moving in unseen physical or virtual environments to carry out human-guided tasks such as locating objects [4, 36, 10, 33], following instructions [19, 13, 1, 31], or reaching a point in space [49, 45]. State-of-the-art methods in this field use Reinforcement Learning (RL)  [47, 45, 20] and Large-Language Model (LLM) based schemes [39, 10] to achieve superior task performance, and these policies also extend to multi-object finding [9, 44]. Topological graphs (TGs) are a common facet of such techniques and are used to store environmental details for decision making. Embodied instruction following tasks such as REVERIE [31]and R2R [1, 13] utilize TGs extracted from MP3D [8] for navigation. Similarly, tasks like ObjectNav, PointNav, and ImageNav also employ TGs [20, 46, 14, 48, 11, 22, 34] where nodes represent locations or landmarks, and edges define traversable paths. These graphs provide a structured way to plan and execute navigation policies.
@@
A key feature of current embodied navigation paradigms [44, 9, 18, 12] is their reliance on static graphs, where objects within nodes remain stationary. This static assumption is limiting, as real-world environments are inherently dynamic, where users frequently move portable objects from place to place over long periods, such as shifting their phone or wallet between rooms throughout the day. Studies on robots in collaborative environments  [29, 2, 16] suggest that human object placements follow structured patterns shaped by routines and habits, introducing a degree of entropy in object locations over time. Despite the important applicability of non-stationary environments, existing navigation methods fail to model such temporal variability, limiting their applicability for real-world deployment.
Figure 1: Transit-Aware Strategy (TAS) We introduce TAS to tackle embodied navigation in non-stationary environments. In this figure, an embodied agent is tasked with finding a wallet that changes positions in the environment between 9:$0$ and 9:20. $P_{\text{wallet}}$ represents the transit route of the wallet. TAS makes agents “transit-aware” by modeling object transit information to augment the agent’s navigation policy for temporally-informed decision making.
In the ObjectNav task [18, 3], for instance, SotA navigation approaches all involve finding stationary targets using RL [47, 45, 3, 41] and LLM-based zero-shot methods [18, 39, 6]. RL policies for multi-object settings [44, 9] have also been proposed, but these are also based on static environments. In contrast, navigating to non-stationary targets that move as the agent moves necessitates the temporal modeling of object transit to guide the agent.
For example, OVRL-V2 [47] uses CLIP text embeddings and ViT-processed RGB images, while DD-PPO [45] similarly relies on GPS readings and RGB-D sensors; both use only static observations of the scene. Their reward functions similarly assume target stationarity, with OVRL-V2 and DD-PPO penalizing absolute distance changes to the goal. Using these approaches directly in non-stationary environments shows poor adaptability because the distance to a target fluctuates as the agent moves, producing a high-variance signal that hinders policy convergence.
@@
For real-world deployment, these approaches must generalize, i.e., maintain performance in dynamic scenes with shifting targets. Prior embodied work describes generalizability in terms of measuring an agent’s capacity to adapt to novel settings, but this usually refers to improved performance in unseen environments or unique synthesized guidance language [50, 25, 23, 17]. Little to no work that defines generalizability in the context of adapting to dynamic settings with non-stationary targets.
Main Results. To address these issues, we present a Transit-Aware Strategy (TAS) that informs embodied navigation policies with temporal “awareness” on target positions (Fig. 1). Rather than naively chasing target objects, TAS augmented agents learn to intercept target routes for better convergence. To evaluate TAS, we introduce Dynamic Object Maps (DOMs), a dynamic variant of topological graphs incorporating structured object transitions. Figure 2 illustrates this concept. DOM node attributes change over time, i.e. target objects move from node to node at each timestep, while the edges remain fixed.A successful agent must synchronize its path with these object routines, which requires learning and anticipating object movement patterns during exploration. This ability reflects an agent’s generalization to non-stationary environments.
Figure 2: Dynamic Object Maps (DOMs): We introduce dynamism to topological graphs via portable targets following human-habit inspired routes. In this representative figure, the watch (in red) moves from node N1 at T=10:$0$ AM to node N5 at T=7:45 PM, while the wallet (in orange) moves from node N5 at T=10:15 AM to node N1 at T=7:45 PM. We benchmark navigation agents under different object transition scenarios and present a Transit-Aware Strategy (TAS) to improve their performance.
Our work seeks to redefine success as defined in traditional embodied navigation tasks, by emphasizing not just on where to navigate but also when to get there i.e. right place, right time. Our contributions are summarized as follows:
  • Transit-Aware Strategy (TAS): We propose TAS, a novel strategy to enhance the capabilities of embodied agents to find non-stationary targets. TAS equips an agent with object transit routes, enabling them to synchronize their paths with moving targets. We develop two TAS-enhanced agents TAS-RL and TAS-LLM, that modify existing RL [45] and LLM [18] policies with temporal transit cues.
@@
  • Dynamic Object Maps (DOMs): We propose an algorithm to introduce structured object transitions on static topological graphs, transforming them into Dynamic Object Maps or DOMs. Inspired by human habits, we define three distinct object transition scenarios to govern the movement of portable objects in a scene, providing varying levels of entropy. DOMs apply to standard TGs from datasets like Matterport3D (MP3D) [8] and HM3D [32].
  • Generalizability Benchmark: We evaluate $6$ different navigation agents on a multi-object finding task across static and DOM environments. Our TAS-RL and TAS-LLM agents show an average improved Success Rate (SR of $22.9\%$ and generalize better by $49.8\%$ in terms of Relative Change in Success (RCS)[17] over non-TAS counterparts.
@@ 2 Related Works
Traditionally, navigation in dynamic environments has focused on low-level planning tasks, such as crowd avoidance [37, 27, 5] and socially aware navigation [7, 26], where agents continuously maneuver to avoid obstacles while minimizing trajectory length or time. In contrast, embodied navigation tasks employing high-level planning on graphs are predominantly studied in static environments with stationary targets, with limited research on handling dynamic targets  [13, 38]. In our work, we introduce dynamism on a graph-level, where agents must reason over long-term spatio-temporal changes of objects for navigation decisions. While recent simulators like Habitat 3.0 [30] include object rearrangement tasks, these primarily focus on agents modifying the environment and not finding non-stationary targets on a graph.
@@
Rudra et. al. [36] define small portable objects around the house and propose a contextual bandit scheme that aims to learn the likelihood of finding an object at various waypoints. In their case, however, object locations are shuffled only after each episode, meaning it finally boils down to finding stationary targets at each timestep. In contrast, we tackle a truly dynamic case, where objects move even during the episode. This definition adds a layer of complexity as the embodied agent must now plan to navigate towards a shifting target, requiring an explicit modeling of the object’s transit.
Kurenkov et. al. [21] introduced dynamizing household environments and experimented with scene graph memory to predict object locations. They also performed target object-finding experiments, but the target was stationary as the agent moved. In our work, the objects move as the agent moves, and we deal with topological graphs, not scene graphs. Wang et. al in [43] used LLM-generated human activities to dynamize a topological graph of a household environment and performed ObjectNav experiments in these environments. In contrast, our work presents a transit-aware strategy to improve navigation performance in non-stationary environments.
3 Method: Finding Non-Stationary Targets
In this section, we first our approach to developing an environment with non-stationary targets with human-habit inspired transit scenarios. We then present our strategy to improve embodied navigation performance on these environments by integrating agents with transit-awareness.
3.1 Defining Transit Scenarios $\Lambda$:
Human habit formation via object placement can be attributed to cognitive offloading, where individuals rely on the environment to reduce memory demands [35, 40]. This reliance on the environment influences the spatial distribution of objects. Frequently used objects (like keys or phones) might move more flexibly across rooms (high entropy), while function-specific objects (like dumbbells or toothbrush) might be more constrained (low entropy). We use this structured transition behavior to define an object’s transition entropy, which we use to formulate $\Lambda$. We consider three object transit scenarios at decreasing levels of transit entropy, Random, Semi-Routine, and Fully-Routine. These are outlined in Table 1 below.
Scenario ($\mathbf{\Lambda}$) Fixed Rooms Fixed Paths Entropy
Static N/A None
Random High
Semi-Routine Medium
Fully-Routine Low
Figure 3: Object Transition: Portable objects move around the scene at various timesteps in accordance to their natural rooms (Table I in Appendix) and transit scenarios (Table 1). Here, a mug is placed on a kitchen node $N_{3}$ at timestep $T=53$, but moves to a bedroom node $N_{5}$ at timestep $T=55$ via node $N_{4}$. If an agent reaches the kitchen after $T=53$ (or the bedroom before $T=55$), it would fail to see the mug. Multiple objects could also be at the same node (hat and mug in $N4$).
3.2 Dynamic Object Maps (DOMs):
We define Dynamic Object Maps (DOMs) as node-attributed dynamic graphs where edge values remain constant, but the node attributes (objects) evolve over time. DOMs extend attributed graphs with temporal object dynamics. Formally, let $G=(V,E)$ be a topological graph structure with nodes $V$ representing spatial locations (rooms and objects), and edges $E$ representing traversable paths between the nodes. Let $\mathcal{O}$ represent the total set of available portable objects. Then, at each timestep $t$, $v\in V$ is associated with a set of portable target objects $O_{v}(t)\subseteq\mathcal{O}$. While the graph edges $E$ remain fixed over time, the node attributes or object assignments $O_{v}(t)$ evolve via a time-dependent transit scenario $\Lambda$, to give us an object transition function $O$ as,